# Cascades towards noise-induced transitions on networks revealed using information flows

### v3

@@ 1 Introduction
Multistability is a crucial characteristic observed in various real-world complex systems [1, 2]. It describes the phenomenon in which a system, under the influence of noise, explores different states within its state space at various spatio-temporal scales. Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [3, 4]. Larger-scale instances include multistability in opinion dynamics [5] and ecosystems or climate systems balancing between metastable states [6, 7].
Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [8, 9, 10, 11]. Traditionally, noise is perceived as an unwanted signal, diminishing system effectiveness or signal quality. However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [12, 13].
Here, we focus on a particular subset of dynamical systems where each node’s state in the next time step is governed by the Boltzmann-Gibbs distribution, akin to a local thermal equilibrium perspective. Such models have been applied to describe various behaviors, including neural dynamics [14], opinion dynamics, ferromagnetic spins [15], and organized criminal gang interactions [16].
@@
In physical applications, this transition involves overcoming an energy barrier (fig. 1), akin to frustration, assortativity, or a fitness score. A global ‘temperature’ parameter introduces noise, influencing the likelihood of choosing higher energy states. For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [17, 18, 19, 7]. Our method uncovers the network percolation process facilitating endogenously generated, noise-induced transitions, offering a different perspective from traditional approaches. This computational method requires only access to system state probabilities over time, making it broadly applicable.
@@ 2 Methods
Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [20] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$) $I(s_{i}^{\tau}:S^{\tau+t})$. The time lag $t$ is used to analyze two key features of information flows of a system: the area under the curve (AUC) of short-term information, and sustained level of long term information.
The contribution of a node to the dynamics of the system will differ depending on the network connectivity of a node (fig. A.8) [21, 22]. The total amount of fluctuations shared between the node’s current state and the system’s short-term future trajectory is computed as the integrated mutual information
@@
Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [21]. It reflects how much of the node state is in the “working memory” of the system.
@@ 4 Discussion
Integrated mutual information captured how short-lived correlations are passed on from the initator nodes. In the stable regime (close to the ground state) low degree nodes drive the system dynamics. Low degree nodes destabilize the system, pushing the system closer to the tipping point. In most cases, the initiator nodes will fail in propagating the noise to their neighbors. On rare occasions, however, the cascade is propagated progressively from low degree, to higher and higher degree. A similar domino mechanism was recently found in climate science [19, 7]. Wunderling and colleagues provided a simplified model of the climate system, analyzing how various components contribute to the stability of the climate. They found that interactions generally stabilize the system dynamics. If, however, a metastable transitions was initialized, noise was propagated through a similar mechanism as found here. That is, an “initializer” node propagated noise through the system which created a domino effect that percolated through the system. The results from this study mirrors these conclusions and provides a model-free language to express these domino effects.
@@ 6 Limitations
Information flows conditioned on a partition is a form of conditional mutual information [23]. Prior results showed that conditional information produces synergy, i.e. information that is only present in the joint of all variables but cannot be found in any of the subset of each variable. Unfortunately, there is no generally agreed upon definition on how to measure synergy [24, 25] and different estimates exist that may over or underestimate the synergetic effects. By partitioning one can create synergy as for a given partition each spin has some additional information about the other spins. For example, by taking the states such that $\langle S\rangle=0.1$, each spin “knows” that the average of the system equals 0.1. This creates shared information among the spins. Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[26]. Using this approach, no synergy was measured that affected the outcome of this study. However, it should be emphasized that synergetic effects may influence the causal interpretation of the approach presented here.
@@ A.1 Background, scope & innovation
Noise induced transitions produces produces metastable behavior that is fundamental for the functioning of complex dynamical systems. For example, in neural systems, the presence of noise increases information processing. Similarly, the relation between glacial ice ages and earth eccentricity has been shown to have a strong correlation. Metastability manifests itself by means of noise that can be of two kinds [27]. External noise originates from events outside the internal system dynamics [28, 29]. Examples include the influence of climate effects, population growth or a random noise source on a transmission line. External noise is commonly modeled by replacing an external control or order parameter by a stochastic process. Internal noise, in contrast, is inherent to the system itself and is caused by random interactions of elements of the system, e.g. individuals in a population, or molecules in chemical processes. Both types of noise can generate transitions between one metastable state and another. In this paper, the metastable behavior is studied of internal noise in complex dynamical networks governed by the kinetic Ising dynamics.
@@
The ubiquity of multistability in complex systems calls for a general framework to understand how metastable transitions occur. The diversity of complex systems can be captured by an interaction networks that dynamically evolves over time. These dynamics can be seen as a distributive network of computational units, where each unit or element of the interaction network changes it state based on the input it gets from its local neighborhood. Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [30, 31, 32]. Instead of describing the dynamics of the system in terms of their domain knowledge such as voltage over distance, disease spreading rate, or climate conditions, one can understand the dynamics in terms of the information dynamics. In particular, the field of information dynamics is concerned with describing the system behavior along its capacity to store information, transmit information, and modify information. By abstracting away the domain details of a system and recasting the dynamics in terms of how the system computes its next state, one can capture the intrinsic computation a system performs. The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [33].
@@
Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [20]. Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [13, 34].
@@
Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [35, 36, 37]. It often centers around an observation that the system’s ability to absorb noise reduces prior to the system going through a critical point. This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [38]. However, these methods traditionally use some form of control parameter driving the system towards or away from a critical point. Most real-world systems lack such an explicit control parameter and require different methods. Furthermore, detecting a tipping point does not necessarily lead to further understanding how the tipping point was created. For example, for a finite size Ising model, the system produces bistable behavior. As one increases the noise parameter, the bistable behavior disappears. The increase in noise effectively changes the energy landscape, but little information is gained as to how initially the metastable behavior emerged.
@@
In this work, a novel approach using information theory is explored to study metastable behavior. The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [31]. Two information features are introduced. Integrated mutual information measure predictive information of a node on the future of the system. Asymptotic information measures the long timescale memory capacity of a node. These measures differ from previous information methods such as transfer entropy [39], conditional mutual information under causal intervention [40], causation entropy [41], and time-delayed variants [42] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable. Here, instead, we aim to understand how the elements in the system contribute to the macroscopic properties of the system. It is important to emphasize that information flows are not directly comparable to causal flows [23]. A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [31]. In this sense, this work is similar to predictive information [43] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
@@ A.3 Information flow on complex networks
Informally, the information flows measures the statistical coherence between two random variables $X$ and $Y$ over time such that the present information in $Y$ cannot be explained by the past of $Y$ but rather by the past of $X$. Estimating information flow is inherently difficult due to the presence of confounding which potential traps the interpretation in the “correlation does not equal causation”. Under some context, however, information flow can be interpreted as causal [21]. Let $S=\{s_{1},\dots,s_{n}\}$ be a random process, and $S^{t}$ represent the state of the random process at some time $t$. The information present in $S$ is given as the Shannon entropy
@@
Prior results showed that the nodes with the highest causal importance are those nodes that have the highest information flow (i.e. maximize 10) [21]. Intuitively, the nodes for which the future system “remembers” information from a node in the past, is the one that “drives” the system dynamics. Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [21]
@@
In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [21]. However in general information flows are difficult to equate to causal flows [31, 23]. Here, the local information flows are computed by considering the integrated mutual information conditioned on part of the entire state space. This allows for mapping the local information flows between nodes and the system over time, but does not guarantee that the measured information flows are directly causal. The main reason being that having predictive power about the future, could be completely caused by the partitioning. In [21] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
In addition, in [21] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered. Applying this approach under a state partition $I(S^{\tau}:s_{i}^{\tau+t}|\langle S\rangle)$causes a violation of the data processing as information may flow from a node at a particular $t=t_{1}$ and then flow back to the node at $t=t2,t_{2}>t_{1}$. In order to simplify the interpretation of the information flows and keep the data processing inequality, the reverse $I(S^{t+\tau}:s_{i}^{\tau}|\langle S\rangle)$ was computed in the present study.
@@ A.4 Noise matching procedure
The Boltzmann-Gibbs distribution is parameterized by noise factor $\beta=\frac{1}{kT}$ where $T$ is the temperature and $k$ is the Boltzmann constant. For high $\beta$ values metastable behavior occurs in the kinetic Ising model. The temperature was chosen such that the statistical complexity [44] was maximized. The statistical complexity $C$ is computed as
@@
A typical statistical complexity curve is seen in fig. A.6. The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [45].

### v4

# ./main.bib

@@ 1 Introduction
Multistability is a crucial characteristic observed in various real-world complex systems [Ladyman2013, vanNes2016]. It describes the phenomenon in which a system, under the influence of noise, explores different states within its state space at various spatio-temporal scales. Examples range from the brain’s ability to operate in one cognitive mode for an extended period before rapidly changing its state [Kandel2000, Fries2015]. Larger-scale instances include multistability in opinion dynamics [Galam2020] and ecosystems or climate systems balancing between metastable states [1986, Wunderling2021].
Evidence increasingly suggests that noise fluctuations play a fundamental role in transitions between attractor states [Beggs2012, Mitchell1993, Mitchella, Forgoston2018a]. Traditionally, noise is perceived as an unwanted signal, diminishing system effectiveness or signal quality. However, in nonlinear systems, such as interconnected neurons in the brain, noise can play a surprising role by enabling the exploration of larger parts of the state space, allowing the system to escape local minima [Czaplicka2013, Nicolis2016].
Here, we focus on a particular subset of dynamical systems where each node’s state in the next time step is governed by the Boltzmann-Gibbs distribution, akin to a local thermal equilibrium perspective. Such models have been applied to describe various behaviors, including neural dynamics [Hopfield1982b], opinion dynamics, ferromagnetic spins [Glauber1963], and organized criminal gang interactions [DOrsogna2015a].
@@
In physical applications, this transition involves overcoming an energy barrier (fig. 1), akin to frustration, assortativity, or a fitness score. A global ‘temperature’ parameter introduces noise, influencing the likelihood of choosing higher energy states. For networked dynamical system, the structure itself significantly influences systemic transitions, making them more likely than in systems without network effects [Harush2017a, Gao2016, Wunderling2020, Wunderling2021]. Our method uncovers the network percolation process facilitating endogenously generated, noise-induced transitions, offering a different perspective from traditional approaches. This computational method requires only access to system state probabilities over time, making it broadly applicable.
@@ 2 Methods
Fluctuations and their correlations at time $\tau$ are captured using Shannon’s mutual information [Cover2005] shared between a node’s state ($s_{i}^{t}$) at time $t$ and the entire future system state ($S^{t+\tau}$) $I(s_{i}^{\tau}:S^{\tau+t})$. The time lag $t$ is used to analyze two key features of information flows of a system: the area under the curve (AUC) of short-term information, and sustained level of long term information.
The contribution of a node to the dynamics of the system will differ depending on the network connectivity of a node (fig. A.8) [vanElteren2022, Quax2013]. The total amount of fluctuations shared between the node’s current state and the system’s short-term future trajectory is computed as the integrated mutual information
@@
Intuitively, $\mu(s_{i})$ represents a combination of the intensity and duration of the short-term fluctuations on the (transient) system dynamics [vanElteren2022]. It reflects how much of the node state is in the “working memory” of the system.
@@ 4 Discussion
Integrated mutual information captured how short-lived correlations are passed on from the initator nodes. In the stable regime (close to the ground state) low degree nodes drive the system dynamics. Low degree nodes destabilize the system, pushing the system closer to the tipping point. In most cases, the initiator nodes will fail in propagating the noise to their neighbors. On rare occasions, however, the cascade is propagated progressively from low degree, to higher and higher degree. A similar domino mechanism was recently found in climate science [Wunderling2020, Wunderling2021]. Wunderling and colleagues provided a simplified model of the climate system, analyzing how various components contribute to the stability of the climate. They found that interactions generally stabilize the system dynamics. If, however, a metastable transitions was initialized, noise was propagated through a similar mechanism as found here. That is, an “initializer” node propagated noise through the system which created a domino effect that percolated through the system. The results from this study mirrors these conclusions and provides a model-free language to express these domino effects.
@@ 6 Limitations
Information flows conditioned on a partition is a form of conditional mutual information [James2016]. Prior results showed that conditional information produces synergy, i.e. information that is only present in the joint of all variables but cannot be found in any of the subset of each variable. Unfortunately, there is no generally agreed upon definition on how to measure synergy [Beer2015, Kolchinsky2022] and different estimates exist that may over or underestimate the synergetic effects. By partitioning one can create synergy as for a given partition each spin has some additional information about the other spins. For example, by taking the states such that $\langle S\rangle=0.1$, each spin “knows” that the average of the system equals 0.1. This creates shared information among the spins. Analyses were performed to estimate synergy using the redundancy estimation $I_{min}$[Williams2010]. Using this approach, no synergy was measured that affected the outcome of this study. However, it should be emphasized that synergetic effects may influence the causal interpretation of the approach presented here.
@@ A.1 Background, scope & innovation
Noise induced transitions produces produces metastable behavior that is fundamental for the functioning of complex dynamical systems. For example, in neural systems, the presence of noise increases information processing. Similarly, the relation between glacial ice ages and earth eccentricity has been shown to have a strong correlation. Metastability manifests itself by means of noise that can be of two kinds [Forgoston2018]. External noise originates from events outside the internal system dynamics [Calim2021, Czaplicka2013a]. Examples include the influence of climate effects, population growth or a random noise source on a transmission line. External noise is commonly modeled by replacing an external control or order parameter by a stochastic process. Internal noise, in contrast, is inherent to the system itself and is caused by random interactions of elements of the system, e.g. individuals in a population, or molecules in chemical processes. Both types of noise can generate transitions between one metastable state and another. In this paper, the metastable behavior is studied of internal noise in complex dynamical networks governed by the kinetic Ising dynamics.
@@
The ubiquity of multistability in complex systems calls for a general framework to understand how metastable transitions occur. The diversity of complex systems can be captured by an interaction networks that dynamically evolves over time. These dynamics can be seen as a distributive network of computational units, where each unit or element of the interaction network changes it state based on the input it gets from its local neighborhood. Lizier proposed that these proposed that the dynamic interaction of complex systems can be understood by their local information processing [Lizier2008, Lizier2013, Lizier2018]. Instead of describing the dynamics of the system in terms of their domain knowledge such as voltage over distance, disease spreading rate, or climate conditions, one can understand the dynamics in terms of the information dynamics. In particular, the field of information dynamics is concerned with describing the system behavior along its capacity to store information, transmit information, and modify information. By abstracting away the domain details of a system and recasting the dynamics in terms of how the system computes its next state, one can capture the intrinsic computation a system performs. The system behavior is encoded in terms of probability, and the relationship among these variables are explored using the language of information theory [Quax2017].
@@
Information theory offers profound benefits over traditional methods used in meta-stability analysis as the methods developed are model-free, can capture non-linear relationships, can be used for both discrete and continuous variables, and can be estimated directly from data [Cover2005]. Shannon information measures such as mutual information and as well as Fisher information can be used to study how much information the system dynamics shares with the control parameter [Nicolis2016, Lizier2010].
@@
Past research on information flows and metastable transitions focuses on methods to detect the onset of a tipping point [Scheffer2009, Prokopenko2011, Scheffer2001]. It often centers around an observation that the system’s ability to absorb noise reduces prior to the system going through a critical point. This critical slowing down, can be captured as a statistical signature where the Fisher information peaks [Eason2014]. However, these methods traditionally use some form of control parameter driving the system towards or away from a critical point. Most real-world systems lack such an explicit control parameter and require different methods. Furthermore, detecting a tipping point does not necessarily lead to further understanding how the tipping point was created. For example, for a finite size Ising model, the system produces bistable behavior. As one increases the noise parameter, the bistable behavior disappears. The increase in noise effectively changes the energy landscape, but little information is gained as to how initially the metastable behavior emerged.
@@
In this work, a novel approach using information theory is explored to study metastable behavior. The statistical coherence between parts of the system are quantified by the the capability of individual nodes to predict the future behavior of the system [Lizier2013]. Two information features are introduced. Integrated mutual information measure predictive information of a node on the future of the system. Asymptotic information measures the long timescale memory capacity of a node. These measures differ from previous information methods such as transfer entropy [Schreiber], conditional mutual information under causal intervention [Ay2008], causation entropy [Runge2019], and time-delayed variants [Li2008] in that these methods are used to infer the transfer of information between sets of nodes by possible correcting for a third variable. Here, instead, we aim to understand how the elements in the system contribute to the macroscopic properties of the system. It is important to emphasize that information flows are not directly comparable to causal flows [James2016]. A rule of thumb is that causal flows focus on micro-level dynamics ($X$ causes $Y$), whereas information flows focus on the predictive aspects, a holistic view of emergent structures [Lizier2013]. In this sense, this work is similar to predictive information [Bialek1999] where predictive information of some system $S$ is projected onto its consistent elements $s_{i}\in S$ and computed as a function of time $t$.
@@ A.3 Information flow on complex networks
Informally, the information flows measures the statistical coherence between two random variables $X$ and $Y$ over time such that the present information in $Y$ cannot be explained by the past of $Y$ but rather by the past of $X$. Estimating information flow is inherently difficult due to the presence of confounding which potential traps the interpretation in the “correlation does not equal causation”. Under some context, however, information flow can be interpreted as causal [vanElteren2022]. Let $S=\{s_{1},\dots,s_{n}\}$ be a random process, and $S^{t}$ represent the state of the random process at some time $t$. The information present in $S$ is given as the Shannon entropy
@@
Prior results showed that the nodes with the highest causal importance are those nodes that have the highest information flow (i.e. maximize 10) [vanElteren2022]. Intuitively, the nodes for which the future system “remembers” information from a node in the past, is the one that “drives” the system dynamics. Formally, these driver nodes can be identified by computing the total information flow between $S^{t}$ and $s_{i}$ can be captured with the integrated mutual information [vanElteren2022]
@@
In some context, the nodes that maximizes the (11) are those nodes that have the highest causal influence in the system [vanElteren2022]. However in general information flows are difficult to equate to causal flows [Lizier2013, James2016]. Here, the local information flows are computed by considering the integrated mutual information conditioned on part of the entire state space. This allows for mapping the local information flows between nodes and the system over time, but does not guarantee that the measured information flows are directly causal. The main reason being that having predictive power about the future, could be completely caused by the partitioning. In [vanElteren2022] the correlation measured considered all possible states, and the measures were directly related to a causal effect.
In addition, in [vanElteren2022] the shared information between the system with a node shifted over time ($I(S^{\tau}:s_{i}^{\tau+t})$) was considered. Applying this approach under a state partition $I(S^{\tau}:s_{i}^{\tau+t}|\langle S\rangle)$causes a violation of the data processing as information may flow from a node at a particular $t=t_{1}$ and then flow back to the node at $t=t2,t_{2}>t_{1}$. In order to simplify the interpretation of the information flows and keep the data processing inequality, the reverse $I(S^{t+\tau}:s_{i}^{\tau}|\langle S\rangle)$ was computed in the present study.
@@ A.4 Noise matching procedure
The Boltzmann-Gibbs distribution is parameterized by noise factor $\beta=\frac{1}{kT}$ where $T$ is the temperature and $k$ is the Boltzmann constant. For high $\beta$ values metastable behavior occurs in the kinetic Ising model. The temperature was chosen such that the statistical complexity [Lopez-Ruiz1995a] was maximized. The statistical complexity $C$ is computed as
@@
A typical statistical complexity curve is seen in fig. A.6. The noise parameter $\beta$ is set such that it maximizes the statistical complexity using numerical optimization (COBYLA method in scipy’s optimize.minimize module) [Virtanen2020].